# BlackboxBench: A Comprehensive Benchmark of Black-box Adversarial Attacks

### v1

@@ IV-G3 Intra/Cross-training-strategy transferability
To comprehensively assess the impact of different training strategies, in addition to normally trained ResNet50, ConvNeXt-B, Swin-B the same as mentioned in Sec. IV-E, we also exploit their adversarially trained counterparts, denoted as ResNet50(AT)[58], ConvNeXt-B(AT)[59] and Swin-B(AT)[59]. This study considers all possible combinations of these models as surrogate-target pairs. For each pair, we subject them to the untargeted, $\ell_{\infty}$ attack on the ImageNet and present the transferability between them in Fig. 21. We can see that, in general, the transferability between regular models and AT models is extremely low. For example, in DI2-FGSM, the model with highest transferability to ResNet50(AT) is ConvNeXt-B(AT) instead of ResNet50, the model with highest transferability to ConvNeXt-B is Swin-B instead of ConvNeXt-B(AT). That is to say, strong transferability mostly occurs within the same training strategies, but not between different strategies.
@@
We gain intuition on the cause of this phenomenon by analyzing the spatial frequency features utilized by models to make predictions, via our analysis tool—FSM. The FSMs of these six models are shown in Fig. 21. Firstly, we find that FSMs of normal models consistently display uniform attention distribution across the frequency spectrum, with certain frequency components (e.g., very low frequency) universally disregarded. We also observe the FSM similarity among AT models: their attention spreads over the whole frequency spectrum and predominantly focuses on the frequencies ignored by the normal counterparts. As per Eq. (10), FSM is the Discrete Fourier Transform (DFT) of model gradient w.r.t input. Given that adversarial perturbations are largely parallel to the gradient, the magnitude of the Fourier spectrum of adversarial perturbation crafted on a model aligns with the magnitude of the model’s FSM, $i.e.$, the perturbation span frequency ranges which the model is sensitive to. Due to the similarity in frequency attention maps among models with the same training strategy, the crafted perturbations also maintain similarity. The transferability of adversarial perturbations between models arises from shared non-robust features. Consequently, perturbations transfer effectively between models with the same training strategy, yet exhibit reduced transferability between models employing different training strategies. Note that all models transfer well to ResNet50 might be attributed to their coverage of a broader frequency spectrum compared to ResNet50.
V Conclusion
The BlackboxBench benchmark, featuring a modular, unified and extensible codebase, can contribute to validating the truly effective ideas within the domain of black-box adversarial attacks. Our implementation encompasses 25 query-based attacks (13 score-based and 12 query-based), along with 30 transfer-based attacks (23 single and 7 composed). These efforts resulted in 14,106 evaluations covering a wide range of attack scenarios. Our aim with BlackboxBench is to facilitate the research community, especially in rapidly comparing with existing literature during the design of new attacks, providing a sound grasp of the progress in black-box adversarial attacks over time, and assessing defenses against adversarial threats at inference time. Furthermore, we demonstrated that our tools are valuable for analyzing various factors related to the vulnerability of DNNs, such as data, model architecture, attack budget, and defense mechanisms. We believe that the research community needs to gain a deeper understanding of how these factors impact attack effectiveness. In this endeavor, BlackboxBench can play a pivotal role.
Appendix A Implementation details on experiments
A-A Pretrained model download links
A-B Experimental setting for Result Overview in Sec. IV-B
@@ A-A Pretrained model download links
In the experiments of query-based attacks, we implemented both untargeted and targeted attacks. For untargeted attacks, we assign the groud-truth label provided by the validation set for each clean example. In this case, one attack is successful if the predicted class of the adversarial example is different from the ground-truth label. The maximum number of queries is set to 10,000 for all untargeted attacks. For targeted attacks, we also assign the groud-truth label as untargeted attacks for each clean example. Besides, we choose the target label for each clean example according to the logit returned by the target model. There are 5 target types of target labels, which are "random", "least likely", "most likely", "median" and "specified label", for users to choose from. In our evaluation, we choose to assign a target label for each clean example by "median", which is the median label of the logit given by the target model. For all targeted attacks, the maximum number of queries is set to 100,000 since if we still set the number to 10,000, which is the same as untargeted attacks, most attacks will fail to craft even one adversarial example. In Sec. IV-B we only select the attack setting with ImageNet as our dataset and ResNet-50 as our target model. We present the performance of both $\ell_{2}$ norm and $\ell_{\infty}$ norm attacks.
In the experiments of transfer-based attacks, we test all four attack settings on ImageNet. We employ ResNet50 as the surrogate model and encompass all 17 models listed in Tab. III as target models. The reported performances are averaged ASR over all target models. For untargeted attacks, we utilize Cross Entropy loss. The adversarial perburbation is bounded by $\ell_{\infty}=0.03$ with step size $1/255$ and $\ell_{2}=3$ with step size $0.3$ for all attacks. The number of iteration is method-specific. For targeted attacks, we utilize the logit loss. The adversarial perburbation is bounded by $\ell_{2}=0.06$ with step size $0.006$ and $\ell_{2}=6$ with step size $0.6$ for all attacks. The number of iteration varies by method, generally exceeding that of untargeted attacks.
A-C Experimental setting for Effect of Data in Sec. IV-D
@@
A-D Experimental setting for Effect of Model Architecture in Sec. IV-E
@@
A-E Implementation details for Effect of Attack budget in Sec. IV-F
@@
A-F Experimental setting for Effect of Defense in Sec. IV-G
@@
Appendix B Hyper-parameters
The hyper-parameters for score-based attacks, decision-based attacks and transfer-based attacks are specified in Tab. V, Tab. IV and Tab. VI. Due to potential variation in effectiveness across the different surrogate model architectures and datasets, hyper-parameter configurations are meticulously detailed for each scenario. One can reproduce all the evaluation results with these configurations. Additionally, we provide our rules to select the hyper-parameters, as follows:
@@ Appendix B Hyper-parameters
Appendix C Unified algorithms
BlackboxBench unifies decision-based attacks, score-based attacks and transfer-based attacks in Alg.1, Alg.2 and Alg.3 respectively. Any attack process in BlackboxBench could be composed of the functional blocks in this algorithm. The algorithms are described in Sec. III-C.
C-A Decision-based attack algorithm
Input: Target model $f$, a benign example $\boldsymbol{x}$, label $y$;
Hyper-parameter: Maximum perturbation $\epsilon$, step size $\alpha$, max query numbers $N$;
Function specific to attacks: Adversarial example initialization function $\mathcal{I}$, perturbation generation function $\mathcal{G}$;
1: Initialize: $\boldsymbol{x}_{0}^{*}\leftarrow\mathcal{I}(\boldsymbol{x})$;
2:while $t<N$ and ${\left\|\boldsymbol{g}_{t}\right\|_{p}}>\epsilon$ do
3:    # Perturbation generation
4:    Generate the perturbation $\boldsymbol{g}_{t+1}$ using $\mathcal{G}\left(\boldsymbol{x}_{t}^{*}\right)$;
5:    # Acceptance check
6:    if $\boldsymbol{g}_{t+1}$ satisfies acceptance check then
7:       # Update step
8:       Generate adversarial example $\boldsymbol{x}_{t+1}^{*}$ using an update step based on $\boldsymbol{g}_{t+1}\text{ and }\alpha$;
9:    end if
10:end while
11:return  $\boldsymbol{x}^{*}=\boldsymbol{x}_{N}^{*}$.
@@ C-A Decision-based attack algorithm
C-B Score-based attack algorithm
Input: Target model $f$, a benign example $\boldsymbol{x}$, label $y$;
Hyper-parameter: Maximum perturbation $\epsilon$, step size $\alpha$, max query numbers $N$;
Function specific to attacks: Adversarial example initialization function $\mathcal{I}$, perturbation generation function $\mathcal{G}$;
1: Initialize: $\boldsymbol{x}_{0}^{*}\leftarrow\mathcal{I}(\boldsymbol{x})$;
2:while $t<N$ and $\boldsymbol{x}_{0}^{*}$ is not adversarial do
3:    # Perturbation generation
4:    Generate the perturbation $\boldsymbol{g}_{t+1}$ using $\mathcal{G}\left(\boldsymbol{x}_{t}^{*}\right)$;
5:    # Acceptance check
6:    if $\boldsymbol{g}_{t+1}$ satisfies acceptance check then
7:       # Update step
8:       Generate adversarial example $\boldsymbol{x}_{t+1}^{*}$ using an update step based on $\boldsymbol{g}_{t+1},\alpha,\epsilon$
9:    end if
10:end while
11:return  $\boldsymbol{x}^{*}=\boldsymbol{x}_{N}^{*}$.
@@ C-B Score-based attack algorithm
C-C Transfer-based attack algorithm
@@
Appendix D Description of black-box adversarial attacks
@@
D-A Description of decision-based attacks
@@
D-B Description of score-based attacks
@@
D-C Description of transfer-based attacks
  • Iterative Fast Gradient Sign Method (I-FGSM) [20] I-FGSM is the first work to optimize function5 multiple times with a small step size and iteratively update adversarial examples. Compared with the one-step attack FGSM[107], I-FGSM performs better in the white-box setting and worse in the black-box setting.
@@ D-C Description of transfer-based attacks
  • Projected Gradient Descent (PGD)[21] PGD is an iterative version of FGSM[107] but starts at a random perturbation in the $\ell_{p}$-ball at each run.
@@
  • Translation-invariant Fast Gradient Sign Method (TI-FGSM)[22]: As adversarial examples generated against a white-box model are highly correlated with the discriminative region of the white-box model at the given input point, the tranferability to models with different discriminative regions is reduced. TI-FGSM shifts images by translation operation to generate adversarial examples less sensitive to the discriminative regions of the white-box model.
@@
Model perspective
@@
Composite Attack
  • BlackboxBench provides 7 composite attacks. MI-DI is the composition of MI and DI. MI-DI-TI is the compisition of MI, DI and TI. MI-DI-TI-SI is the composition of MI, DI, TI, and SI. VMI is the composition of VT and MI. VNI is the composition of VT and NI. SI-RAP is the composition of SI and RAP. LGV-GhostNet is the composition of LGV and GhostNet.
Appendix E Description of black-box adversarial defenses
E-A Description of static defenses

### v2

@@ C-C3 Intra/Cross-training-strategy transferability
To comprehensively assess the impact of different training strategies, in addition to normally trained ResNet50, ConvNeXt-B, Swin-B the same as in Sec. IV-D2, we also exploit their adversarially trained counterparts, denoted as ResNet50(AT)[97], ConvNeXt-B(AT)[98] and Swin-B(AT)[98]. This study considers all possible combinations of these models as surrogate-target pairs. For each pair, we subject them to the untargeted, $\ell_{\infty}$ attack on the ImageNet and present the transferability between them in Fig. 25. We can see that, in general, the transferability between regular models and AT models is extremely low. For example, in DI2-FGSM, the model with highest transferability to ResNet50(AT) is ConvNeXt-B(AT) instead of ResNet50, the model with highest transferability to ConvNeXt-B is Swin-B instead of ConvNeXt-B(AT). That is to say, strong transferability mostly occurs within the same training strategies, but not between different strategies.
@@
We gain intuition on the cause of this phenomenon by analyzing the spatial frequency features utilized by models to make predictions, via our analysis tool—FSM. The FSMs of these six models are shown in Fig. 25. Firstly, we find that FSMs of normal models consistently display uniform attention distribution across the frequency spectrum, with certain frequency components (e.g., very low frequency) universally disregarded. We also observe the FSM similarity among AT models: their attention spreads over the whole frequency spectrum and predominantly focuses on the frequencies ignored by the normal counterparts. As per Eq. (I.3), FSM is the Discrete Fourier Transform (DFT) of model gradient w.r.t input. Given that adversarial perturbations are largely parallel to the gradient, the magnitude of the Fourier spectrum of adversarial perturbation crafted on a model aligns with the magnitude of the model’s FSM, $i.e.$, the perturbation span frequency ranges which the model is sensitive to. Due to the similarity in frequency attention maps among models with the same training strategy, the crafted perturbations also maintain similarity. The transferability of adversarial perturbations between models arises from shared non-robust features. Consequently, perturbations transfer effectively between models with the same training strategy, yet exhibit reduced transferability between models employing different training strategies. Note that all models transfer well to ResNet50 might be attributed to their coverage of a broader frequency spectrum compared to ResNet50.
C-D Attack Procedure Analysis
Figure 29: The rest part of Fig. 28
For a more qualitative analysis, we explore what is most salient to the target model when an image is under some attack. Using our analytical tools—FullGrad, we visualize the evolution of model attention maps on the perturbed images along the iteration in transfer-based attacks, or the path traversed by query-based attackers. We are interested in both how successful attacks succeed and how failed attacks fail. The main observation is that, in both transfer-based and query-based attacks, successful adversarial attacks manage to shift the model’s attention from critical, discriminative regions to trivial areas, leading to misclassification, whereas failed attacks cannot divert the model’s focus from the key regions, allowing the model to correctly classify the images despite the attack.
In query-based attacks, we choose an image from the category of squirrel, which is also selected to be successfully attacked by about $80\%$ of all methods. Following the similar presentation of FullGrad in transfer-based attacks, we show the attention maps through the attack process of unsuccessful attacks in Fig. 30 and successful attacks of both decision-based attacks and score-based attacks in Fig. 31 and Fig. 32. We can also observe that in samples where attacks failed, the target model continuously concentrated on the critical regions or nearby regions, which failed to mislead the target model and thus the target model can correctly identify them throughout the attack process. Conversely, during successful attacks, the adversarial images incrementally shift the model’s attention away from these critical regions (from the central part to another part). Eventually, the models are deceived into concentrating on insignificant areas, leading to the misclassification of the adversarial images as the category of a beaver.
D Implementation details on experiments
D-A Pretrained model download links
For evalutions on CIFAR-10 models, we use the following implementations(embedded links):
Models for ImageNet dataset are from following implementations:
The clean accuracies of above models on benign test sets are shown in Tab. IV. On CIFAR-10, they are evaluated on 10,000 test set. On ImageNet, they are evaluated on the 50,000 validation set.
TABLE IV: Clean accuracies of the models adopted in BlackboxBench. Models in blue are treated as surrogate models. AT indicates adversarially trained models. PyramidNet indicates PyramidNet[105] + ShakeDrop[106] + AutoAugment[107]. VGG-19 indicates VGG-19 with batch normalization.
Query CIFAR-10 ResNet-50[83] VGG-19[84] Inception-V3[85] DenseNet-121[108] WRN-28-10 (AT w/ DDPM) [93] WRN-28-10 (AT w/ optimization trick) [94]
93.65% 93.95% 93.74% 94.06% 60.75% 62.80%
ImageNet ResNet-50[83] VGG-19[84] Inception-V3[85] ConvNeXt-S[86] ViT-B/16[87] ResNet-50 (AT-$\ell_{2}$) [97] ResNet-50 ($\ell_{\infty}$)[97] ResNet-50 (FastAT)[95]
76.13% 74.22% 77.29% 83.62% 81.07% 62.83% 54.53% 43.46%
Transfer CIFAR-10 ResNet-50[83] VGG-19[84] Inception-V3[85] DenseNet-BC[108] ResNeXt-29[109] WRN-28-10[110] PyramidNet[66] GDAS[111] WRN-28-10 (AT)[96]
95.30% 93.34% 94.77% 96.68% 96.31% 96.21% 98.44% 97.19% 89.69%
ImageNet ResNet-50[83] VGG-19[84] Inception-V3[85] DenseNet-121[108] ViT-B/16[87] ResNet-152[83] MobileNet-V2[112] SENet-154[113] ResNeXt-101[109]
76.15% 74.24% 77.29% 74.65% 81.07% 78.31% 71.88% 81.32% 79.31%
WRN-101[110] PNASNet[114] MNASNet[115] Swin-B[89] ConvNeXt-B[86] ResNet-50 (AT)[97] Swin-B (AT)[98] ConvNeXt-B (AT)[98]
78.84% 82.90% 73.46% 85.20% 83.80% 64.02% 76.16% 76.02%
D-B Experimental setting for Result Overview
@@
In the experiments of query-based attacks, we implemented both untargeted and targeted attacks. For untargeted attacks, we assign the groud-truth label provided by the validation set for each clean example. In this case, one attack is successful if the predicted class of the adversarial example is different from the ground-truth label. The maximum number of queries is set to 10,000 for all untargeted attacks. For targeted attacks, we also assign the groud-truth label as untargeted attacks for each clean example. Besides, we choose the target label for each clean example according to the logit returned by the target model. There are 5 target types of target labels, which are "random", "least likely", "most likely", "median" and "specified label", for users to choose from. In our evaluation, we choose to assign a target label for each clean example by "median", which is the median label of the logit given by the target model. For all targeted attacks, the maximum number of queries is set to 100,000 since if we still set the number to 10,000, which is the same as untargeted attacks, most attacks will fail to craft even one adversarial example. We only select the attack setting with ImageNet as our dataset and ResNet-50 as our target model. We present the performance of both $\ell_{2}$ norm and $\ell_{\infty}$ norm attacks. The adversarial perturbation is bounded by $\ell_{\infty}=0.03$ and $\ell_{2}=3$ for all attacks.
[…]
D-C Experimental setting for Effect of Data
@@
D-D Experimental setting for Effect of Model Architecture
@@
D-E Implementation details for Effect of Attack budget
@@
D-F Experimental setting for Effect of Defense
@@
E Hyper-parameters
The hyper-parameters for score-based attacks, decision-based attacks and transfer-based attacks are specified in Tab. VI, Tab. V and Tab. VII. Due to potential variation in effectiveness across the different surrogate model architectures and datasets, hyper-parameter configurations are meticulously detailed for each scenario. One can reproduce all the evaluation results with these configurations. Additionally, we provide our rules to select the hyper-parameters, as follows:
@@ E Hyper-parameters
F Unified algorithms
BlackboxBench unifies decision-based attacks, score-based attacks and transfer-based attacks in Alg.1, Alg.2 and Alg.3 respectively. Any attack process in BlackboxBench could be composed of the functional blocks in this algorithm. Based on the unified pipelines, we build our attack module, as illustrated by UML component diagrams in Fig. 34.
Figure 34: UML component diagrams of attack modules for transfer-based attack methods (top), decision-based attack methods (middle), and score-based attack methods (bottom).
F-A Decision-based attack algorithm
Input: Target model $f$, Benign example $\boldsymbol{x}$, Label $y$: The true label of the input example $\boldsymbol{x}$;
Hyper-parameter: Maximum perturbation $\epsilon$, Step size $\alpha$, Max query numbers
Attack-specific Functions: Adversarial example initialization function $\mathcal{I}$, Perturbation generation function $\mathcal{G}$
1:Initialize: Initialize the adversarial example as $\boldsymbol{x}_{0}^{*}\leftarrow\mathcal{I}(\boldsymbol{x})$;
2:while $t<N$ and ${\left\|\boldsymbol{g}_{t}\right\|_{p}}<\epsilon$ do
3:    # Perturbation generation step
4:    Generate the perturbation $\boldsymbol{g}_{t+1}$ using the perturbation generation function $\mathcal{G}\left(\boldsymbol{x}_{t}^{*}\right)$;
5:    # Update adversarial example
6:    Update the adversarial example $\boldsymbol{x}_{t+1}^{*}$ by applying the perturbation $\boldsymbol{g}_{t+1}$ with the step size $\alpha$:
$\boldsymbol{x}_{t+1}^{*}\leftarrow\boldsymbol{x}_{t}^{*}+\alpha\cdot\boldsymbol{g}_{t+1}$
7:    # Acceptance check on the updated adversarial example
8:    if $\boldsymbol{x}_{t+1}^{*}$ satisfies the acceptance criteria (e.g., the adversarial example leads to a misclassification by $f$then
9:       Break the loop
10:    end if
11:end while
12:return  The final adversarial example $\boldsymbol{x}^{*}=\boldsymbol{x}_{N}^{*}$.
@@ F-A Decision-based attack algorithm
F-B Score-based attack algorithm
Input: Target model $f$, Benign example $\boldsymbol{x}$, Label $y$
Hyper-parameter: Maximum perturbation $\epsilon$, Step size $\alpha$, Max query numbers $N$
Attack-specific Functions: Adversarial example initialization function $\mathcal{I}$, Perturbation generation function $\mathcal{G}$
1:Initialize: Initialize the adversarial example as $\boldsymbol{x}_{0}^{*}\leftarrow\mathcal{I}(\boldsymbol{x})$;
2:while $t<N$ and $\boldsymbol{x}_{t}^{*}$ is not adversarial (i.e., $f(\boldsymbol{x}_{t}^{*})$ predicts $y$ with high confidence) do
3:    # Perturbation generation step
4:    Generate the perturbation $\boldsymbol{g}_{t+1}$ using the perturbation generation function $\mathcal{G}\left(\boldsymbol{x}_{t}^{*}\right)$;
5:    # Update adversarial example
6:    Update the adversarial example $\boldsymbol{x}_{t+1}^{*}$ by applying the perturbation $\boldsymbol{g}_{t+1}$ with the step size $\alpha$:
$\boldsymbol{x}_{t+1}^{*}\leftarrow\boldsymbol{x}_{t}^{*}+\alpha\cdot\boldsymbol{g}_{t+1}$
7:    # Check if updated adversarial example is successful
8:    if $f(\boldsymbol{x}_{t+1}^{*})\neq y$ or the confidence of $y$ is sufficiently reduced then
9:       Break the loop
10:    end if
11:end while
12:return  The final adversarial example $\boldsymbol{x}^{*}=\boldsymbol{x}_{t}^{*}$.
@@ F-B Score-based attack algorithm
F-C Transfer-based attack algorithm
@@
G Description of black-box adversarial attacks
@@
G-A Description of decision-based attacks
@@
  • Customized Iteration and Sampling Attack (CISA)[2] CISA) enhances query efficiency by integrating surrogate model information and iterative sampling strategies. CISA addresses the challenges of low query efficiency in decision-based attacks by incorporating a transfer-based phase using a surrogate model. This phase generates intermediate adversarial examples, which are then refined using customized sampling based on noise sensitivity. By utilizing the surrogate model’s gradients to guide dual-direction iterative updates, CISA improves the likelihood of crossing the decision boundary with minimal noise. The approach also adapts step sizes and sampling variance based on previous queries, leading to more efficient noise compression and fewer queries overall.
G-B Description of score-based attacks
@@
  • Subspace Attack[43] Subspace Attack leverages information from a surrogate model by combining it with low-dimensional subspace exploration to enhance the attack’s efficiency. The surrogate model provides an approximate gradient or important directions that guide the selection of the subspace in which the search for adversarial perturbations occurs. By using the surrogate model’s gradient estimates, Subspace Attack refines its search within this reduced space, ensuring that the perturbations are aligned with the target model’s vulnerabilities. This combination of surrogate model information and subspace optimization allows the attack to achieve high success rates with fewer queries.
@@
  • Blackbox Attacks via Surrogate Ensemble Search (BASES)[44] BASES combines surrogate model information with a novel query-efficient optimization technique. The key idea is to use a perturbation machine (PM) that generates adversarial examples by minimizing a weighted loss function over a set of surrogate models. Each surrogate model in the ensemble is assigned a weight, and the weights are iteratively adjusted using feedback from the target model’s output. By leveraging surrogate models, BASES efficiently explores the adversarial space in a low-dimensional search space, significantly reducing the number of queries required while maintaining a high success rate in fooling the target model.
  • Prior-guided Random Gradient-Free (PRGF)[1] PRGF also combines score-based black-box attack techniques with surrogate model information to efficiently estimate the gradient of the target model. The surrogate model provides gradient estimates, which are used to guide the search for adversarial perturbations. PRGF projects random perturbations onto the gradient directions suggested by the surrogate model, refining the estimation of the target model’s gradient based on its output scores. This projection strategy allows PRGF to reduce the number of queries required while maintaining high attack success rates, as the surrogate model helps steer the attack toward the most vulnerable directions.
@@
G-C Description of transfer-based attacks
  • Iterative Fast Gradient Sign Method (I-FGSM) [81] I-FGSM is the first work to optimize the general function of transfer-base attacks multiple times with a small step size and iteratively update adversarial examples. Compared with the one-step attack FGSM[91], I-FGSM performs better in the white-box setting and worse in the black-box setting.
@@ G-C Description of transfer-based attacks
  • Projected Gradient Descent (PGD)[57] PGD is an iterative version of FGSM[91] starting from a random perturbation in the $\ell_{p}$-ball.
@@
  • Penalizing Gradient Norm (PGN)[68] PGN enhances adversarial example transferability by targeting flat local regions of the loss landscape through an efficient approximation of the gradient regularization norm.
@@
  • Translation-invariant Fast Gradient Sign Method (TI-FGSM)[58] As adversarial examples generated against a white-box model are highly correlated with the discriminative region of the white-box model at the given input point, the tranferability to models with different discriminative regions is reduced. TI-FGSM shifts images by translation operation to generate adversarial examples less sensitive to the discriminative regions of the white-box model.
  • Structure Invariant Attack (SIA)[3] Existing input-transformation-based attacks, which apply a single global transformation globally, produce limited diversity in transformed images, thereby restricting their transferability. SIA enhances the diversity of adversarial examples by applying different transformations locally across various parts of an input image without altering the image’s global structure to improve its transferability.
@@
Model perspective: model generation
@@
Model perspective: model fusion
  • Logit ensemble[62] Based on the fact that an adversarial example effective across multiple models may reveal an intrinsic transferable weakness, logit ensemble method attacks multiple models simultaneously by fusing their logit activations.
  • Loss ensemble[62] As an alternative ensemble scheme, loss ensemble method aggregates multiple surrogate models in loss.
  • Longitudinal ensemble[73] By attack only one model from the set of surrogate models rather than using all in each iteration, longitudinal ensemble method significantly reduces computational overhead compared to traditional ensemble approaches, facilitating efficient generation of adversarial examples.
  • Common Weakness Attack (CWA)[76] CWA enhances transferability by exploiting two key properties of model ensembles: the flatness of the loss landscape and proximity to models’ local optima. By leveraging Sharpness Aware Minimization (SAM) and Cosine Similarity Encourager (CSE) to optimize these properties, CWA generates more transferable adversarial examples across different models.
  • Adaptive ensemble attack (AdaEA)[77] To address the limitation of traditional ensemble adversarial attacks that equally fuse outputs from surrogate models, AdaEA adaptively modulates the fusion of model outputs based on their contribution to the adversarial goal and introduces a disparity-reduced filter to align update directions. (It is important to note that the hyperparameters $\eta$ and $\beta$ of AdaEA have a high correlation with the surrogate models used. Care must be taken when setting their values.)
H Description of black-box adversarial defenses
H-A Description of static defenses